OpenStudy (osanseviero):

In a school there are 300 matriculated students (some probability problems)

OpenStudy (osanseviero):

210 of those 300 enter the class. 80% of the students that enter the class get good grades, 15% of the students that dont enter the class get good grades

OpenStudy (osanseviero):

If we chose a random matriculated student, which is the probability of him: 1. Being in the class 2. Don't being in the class and getting good grades 3. Getting good grades

OpenStudy (osanseviero):



OpenStudy (osanseviero):

a. So 210/300=7/10

OpenStudy (anonymous):

id say so too

OpenStudy (osanseviero):

How to do the second?

OpenStudy (anonymous):

first you find how many students that arnt in the class and are getting gud grades so 90 x 15/100 = 27/2 feel sorry for the student that lost a limb :(

OpenStudy (anonymous):

cause apparently in that class there is half of a student :'(

OpenStudy (osanseviero):

wait a sec, please

OpenStudy (osanseviero):

lol

OpenStudy (anonymous):

haha xD well there 90 students not in the class, and only 15% of the 90 pass, so 90 x 15/100

OpenStudy (osanseviero):

90 persons is the 30% of total

OpenStudy (osanseviero):

are you sure?

OpenStudy (anonymous):

and then 27/2 / 300 = 9/200

OpenStudy (anonymous):

umm, i thought i was xD

OpenStudy (osanseviero):

27/2 is the number of persons that got a good grade

OpenStudy (anonymous):

and didnt come to class but now idk

OpenStudy (anonymous):

normally when i got half a person it was a sign that something went wrong but my reasoning makes sense atleast to me lol

OpenStudy (osanseviero):

Can you repeat your reasoning again? (sorry)

OpenStudy (anonymous):

90 students are people that do not come to class, 15% of these students are students that do well in exams. therefore 90 x 15/100 = 27/2

OpenStudy (osanseviero):

Ok, I see now :D

OpenStudy (osanseviero):

And for the third question?

OpenStudy (osanseviero):

We know that 9 people that did not entered got good grades. We need to see the ones that did come in

OpenStudy (osanseviero):

So 80% of 210

OpenStudy (anonymous):

it would be 210 x 80/100 + 27/2

OpenStudy (osanseviero):

168 entered and passed, 9 entered and passed, so 176

OpenStudy (anonymous):

y 9?

OpenStudy (osanseviero):

the 9 did not enter and did pass*

OpenStudy (osanseviero):

Sooo...176...and 44/75 (check please)

OpenStudy (osanseviero):

Probability of entering and getting a good grade

OpenStudy (anonymous):

but it should be 168 + 27/2 because the number of students that didnt come to class and are getting gud grades are 27/2

OpenStudy (anonymous):

so 195/2 students pass and the probability of the teacher choosing one is 195/2 /300

OpenStudy (osanseviero):

Sorry, internet crashed

OpenStudy (osanseviero):

So 181.5/300

OpenStudy (osanseviero):

Weird numbers :/

OpenStudy (osanseviero):

Last one is: Chance of selecting a student that have passed and entered the class

OpenStudy (osanseviero):

7/10 * 8/10 = 56/100 = 38/50= 19/25

OpenStudy (osanseviero):

So,...yes?